# Training robust and generalizable quantum models

### v2

@@ Abstract
Adversarial robustness and generalization are both crucial properties of reliable machine learning models. In this letter, we study these properties in the context of quantum machine learning based on Lipschitz bounds. We derive parameter-dependent Lipschitz bounds for quantum models with trainable encoding, showing that the norm of the data encoding has a crucial impact on the robustness against data perturbations. Further, we derive a bound on the generalization error which explicitly involves the parameters of the data encoding. Our theoretical findings give rise to a practical strategy for training robust and generalizable quantum models by regularizing the Lipschitz bound in the cost. Further, we show that, for fixed and non-trainable encodings, as those frequently employed in quantum machine learning, the Lipschitz bound cannot be influenced by tuning the parameters. Thus, trainable encodings are crucial for systematically adapting robustness and generalization during training. The practical implications of our theoretical findings are illustrated with numerical results.
preprint: APS/123-QED
Introduction.—Robustness of machine learning (ML) models is an increasingly important property, especially when operating on real-world data subject to perturbations. In practice, there are various possible sources of perturbations such as noisy data acquisition or adversarial attacks. The latter are tiny but carefully chosen manipulations of the data, and they can lead to dramatic misclassification in neural networks [1, 2, 3, 4]. It is well-known that adversarial robustness is closely connected to generalization [1, 2, 5, 6], i.e., the ability of a model to extrapolate beyond the training data.
@@ A Lipschitz bound of a model $f$ is any $L>0$ satisfying
for all $x_{1},x_{2}\in\mathcal{D}\subseteq\mathbb{R}^{d}$. By definition, Lipschitz bounds quantify the worst-case output change that can be caused by data perturbations and, thus, they provide a useful measure of adversarial robustness. Therefore, they are a well-established tool for characterizing robustness and generalization properties of ML models [2, 5, 7, 8, 9, 10]. Lipschitz bounds cannot only be used to better understand these two properties, but they also allow one to improve them by regularizing the Lipschitz bound during training [2, 5, 11, 12].
@@
In this letter, we study the interplay of robustness and generalization in quantum machine learning (QML). We consider data re-uploading circuits which generalize classical variational quantum circuits [13, 14, 15] by concatenating a data encoding and a parametrized quantum circuit not only once but repeatedly [16]. This alternation provides substantial improvements on expressivity, leading to a universal quantum classifier even in the single-qubit case [16, 17]. Just as in the classical case, robustness is crucial for quantum models. First, if QML is to provide benefits over classical ML, it is necessary to implement QML circuits which are robust w.r.t. quantum errors occurring due to imperfect hardware in the noisy intermediate-scale quantum (NISQ) era [18]. Questions of robustness of quantum models against such hardware errors have been studied, e.g., in [19, 20]. Lipschitz bounds can be used to study robustness of quantum algorithms against certain types of hardware errors, e.g., coherent control errors [21].
@@
However, robustness against hardware errors is entirely different from and independent of the robustness of a quantum model against data perturbations, which will be the subject of this letter. The latter type of robustness has been studied in the context of quantum adversarial machine learning [22]. Not surprisingly, just like their classical counterparts, quantum models are also vulnerable to adversarial attacks, both when operating based on classical data [23, 24] and quantum data [25, 23, 26]. To mitigate these attacks, it is desirable to design training schemes encouraging adversarial robustness of the resulting quantum model. Existing approaches in this direction include solving an (adversarial) min-max optimization problem during training [23] or adding adversarial examples to the training data set [27]. Our results contribute to quantum adversarial machine learning by studying the robustness of quantum models via Lipschitz bounds, leading to a training scheme for robust quantum models based on Lipschitz bound regularization. We demonstrate the robustness benefits of the proposed regularization with numerical results.
@@
In the present letter, we derive a novel generalization bound which explicitly depends on the parameters of the quantum model, highlighting the role of the data encoding for generalizability. We show with numerical results that regularizing the Lipschitz bound indeed leads to substantial improvements in generalization performance. Finally, given that the derived Lipschitz bound mainly depends on the norm of the data encoding, our results show the importance and benefits of trainable encodings over quantum circuits with a priori fixed encoding as frequently used in variational QML [36, 13, 14, 15, 30, 17].
Quantum models and their Lipschitz bounds.—We consider parametrized unitary operators of the form of
@@
where $\Theta=\{\Theta_{j}\}_{j=1}^{N}$ comprises the set of trainable parameters. The quantum model considered in this letter consists of $U_{\Theta}(x)$ applied to |$\ket{0}$ and followed by a measurement w.r.t. the observable $\mathcal{M}$, i.e.,
@@
The formal derivation can be found in the supplementary material (Section I.B). All results in this letter hold for arbitrary $p$-norms as long as the same $p$ is used for both vector and induced matrix norms.
For a given set of parameters $w_{j}$, (5) allows to compute the Lipschitz bound of the corresponding quantum model. Note that $L_{\Theta}$ depends only on $w_{j}$ but it is independent of $\theta_{j}$. This fact plays an important role for potential benefits of trainable encodings since the parameters $w_{j}$ are not optimized during training for fixed-encoding circuits.
Robustness of quantum models.—Suppose we want to evaluate the quantum model $f_{\Theta}$ at $x$, i.e., we are interested in the value $f_{\Theta}(x)$, but we can only access $f_{\Theta}$ at some perturbed input $x^{\prime}=x+\varepsilon$ with an unknown $\varepsilon$. Such a setup can arise due to various reasons, e.g., $x$ may be the output of some physical process which can only be accessed via noisy sensors. The perturbation $\varepsilon$ may also be the result of an adversarial attack, i.e., a perturbation aiming to cause a misclassification by choosing $\varepsilon$ such that $\lVert f_{\Theta}(x+\varepsilon)-f_{\Theta}(x)\rVert$ is maximized. In either case, to correctly classify $x$ despite the perturbation, we require that $f_{\Theta}(x+\varepsilon)$ is close to $f_{\Theta}(x)$. According to (1), a Lipschitz bound $L$ bounds the maximum possible deviation of $f_{\Theta}(x+\varepsilon)$ from $f_{\Theta}(x)$ as
@@
Regularizing the parameters $w_{j}$ encourages small norms of the data encoding and, thereby, small values of the Lipschitz bound $L_{\Theta}$. The hyperparameter $\lambda>0$ allows for a trade off between the two objectives of a small training loss and robustness/generalization in the cost function. Note that the regularization does not involve the $\theta_{j}$’s since they do not influence the Lipschitz bound (5). Moreover, we do not introduce an explicit dependence of the regularization on $\mathcal{M}$ since we do not optimize over the observable in this work.
@@
We now evaluate our theoretical findings based on the circle classification problem from [37]: within a domain of $\mathcal{X}=[-1,+1]\times[-1,+1]$, a circle with radius $\sqrt{2/\pi}$ is drawn, and all data points inside the circle are labeled with $y=+1$, whereas points outside are labeled with $y=-1$, see the supplementary material (Figure 3). For the quantum model with trainable encoding, we use general SU(2) operators and encode $w_{j}^{\top}x+\theta_{j}$ into the first two rotation angles. We repeat this encoding for each of the considered $3$ qubits, followed by nearest neighbour entangling gates based on CNOTs. Such a layer is then repeated $3$ times. As observable, we use $\mathcal{M}=Z\otimes Z\otimes Z$. The resulting circuit is illustrated in the supplementary material (Figure 5). As norms in the regularized training problem (8), we employ $2$-norms, but we note that exploring different choices is an interesting future direction. For example, a $1$-norm regularization may enforce sparsity and, thereby, simplify implementations on NISQ hardware.
@@
The numerical results for the robustness simulations are shown in Figure 1, where we compare the worst-case test accuracy and Lipschitz bound of three trained models with different regularization parameter $\lambda\in\{0,0.2,0.5\}$. Additionally, the plot shows the accuracy of a quantum model with fixed encoding for the same numbers of qubits and layers (see (11) and the supplementary material (Section III) for details). The worst-case test accuracy of all models is obtained by sampling different noise samples $\varepsilon$ from $[-\bar{\varepsilon},+\bar{\varepsilon}]$ and, thereby, it approximates the accuracy resulting from an adversarial attack norm-bounded by $\bar{\varepsilon}$. As expected, all four models deteriorate with increasing noise level. For zero noise level $\bar{\varepsilon}=0$, the model with the largest regularization parameter $\lambda=0.5$ (and, hence, the smallest Lipschitz bound $L_{\Theta}=3.67$) has a smaller test accuracy than the non-regularized model with $\lambda=0$. This can be explained by a decrease in the training accuracy that is caused by the additional regularization in the cost. For increasing noise levels, however, the enhanced robustness outweighs the loss of training performance and, therefore, the model with $\lambda=0.5$ outperforms the model with $\lambda=0$. The fixed-encoding model achieves comparable performance to the trainable-encoding model with $\lambda=0.5$ for small noise and the worst performance among all models for high noise. These observations can be explained by the reduced expressivity but high Lipschitz bound of the fixed-encoding model. Finally, the model with $\lambda=0.2$ almost always outperforms the model with $\lambda=0$ and, in particular, it yields a higher test accuracy for small noise levels. This can be explained by the improved generalization performance caused by the regularization, an effect we discuss in more detail in the following.
@@
Generalization of quantum models.—The Lipschitz bound (5) not only influences robustness but also has a crucial impact on generalization properties of the quantum model $f_{\Theta}$. Intuitively, a smaller Lipschitz bound implies a smaller variability of $f_{\Theta}$ and, therefore, reduces the risk of overfitting. This intuition is made formal via the following generalization bound.
@@
Figure 2: Results for the generalization simulations. As test set, we draw $10.000$ points uniformly at random and evaluate the trained models with different regularization parameter $\lambda$.
@@
Theorem 1.(informal version) Consider a supervised learning setup with loss $\ell$ and data set $(x_{k},y_{k})\in\mathcal{X}\times\mathcal{Y}$ of size $n$ drawn according to the probability distribution $P$. For the quantum model $f_{\Theta}$ from (4), define the expected risk $\mathcal{R}(f_{\Theta})=\int_{\mathcal{X}\times\mathcal{Y}}\ell(y,f_{\Theta}(x))\mathrm{d}P(x,y)$ and the empirical risk $\mathcal{R}_{n}(f_{\Theta})=\frac{1}{n}\sum_{k=1}^{n}\ell(y_{k},f_{\Theta}(x_{k}))$. The generalization error of $f_{\Theta}$ is bounded as
$\Big{|}\mathcal{R}(f_{\Theta})-\mathcal{R}_{n}(f_{\Theta})\Big{|}\leq$ $\,C_{1}\lVert\mathcal{M}\rVert\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert+\frac{C_{2}}{\sqrt{n}}$ (9)
@@
The detailed version and proof of Theorem 1 are provided in the supplementary material (Section II). Generalization bounds as in (9) quantify the ability of $f_{\Theta}$ to generalize beyond the available data. The bound (9) depends on the data encoding via $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ and on the observable via $\lVert\mathcal{M}\rVert$. In particular, $f_{\Theta}$ achieves a small generalization error if its Lipschitz bound $L_{\Theta}$ is small and the size $n$ of the data set is large. Note, however, the following fundamental trade-off: A too small Lipschitz bound $L_{\Theta}$ may limit the expressivity of $f_{\Theta}$ and, therefore, lead to a high empirical risk $\mathcal{R}_{n}(f_{\Theta})$, in which case the generalization bound (9) is meaningless. In conclusion, Theorem 1 implies a small expected risk $\mathcal{R}(f_{\Theta})$ if $n$ is large, $L_{\Theta}$ is small, and $f_{\Theta}$ has a small empirical risk $\mathcal{R}_{n}(f_{\Theta})$. In contrast to existing generalization bounds [30, 31, 32, 33, 34], the bound (9) is not uniform and explicitly involves the Lipschitz bound (5), i.e., the parameters of the data encoding. Hence, Theorem 1 may allow to overcome the limitations of uniform QML generalization bounds [35].
@@
Benefits of trainable encodings.—A popular class of quantum models is obtained by constructing circuits which alternate between data- and parameter-dependent gates, i.e., replacing $U_{\Theta}(x)$ in (3) by
$U^{\mathrm{f}}_{\varphi}(x)=W(\varphi_{L})V(x)\cdots W(\varphi_{1})V(x),$ (10)
compare [36, 13, 14, 15, 30, 17]. The operators $V,W$ are given by $V(x)=e^{-ix_{D}G_{D}}\cdots e^{-ix_{1}G_{1}}$, $W(\varphi_{j})=e^{-i\varphi_{j,p}S_{p}}\cdots e^{-i\varphi_{j,1}S_{1}}$ for trainable parameters $\varphi_{j}$ and generators $G_{i}=G_{i}^{\dagger}$, $S_{i}=S_{i}^{\dagger}$. The corresponding quantum model is given by
$f^{\mathrm{f}}_{\varphi}(x)=\braket{0}{U^{\mathrm{f}}_{\varphi}(x)^{\dagger}\mathcal{M}U^{\mathrm{f}}_{\varphi}(x)}{0},$ (11)
see the supplementary material (Figure 6).
It is not hard to show that the parametrized quantum circuit $U_{\Theta}(x)$ in (3) generalizes the one in (10). Note that the data encoding of the quantum model $f^{\mathrm{f}}_{\varphi}(x)$ is fixed a priori via the choice of $w_{j}$ and, in particular, it cannot be influenced during training. Therefore, we refer to $f^{\mathrm{f}}_{\varphi}(x)$ as a quantum model with fixed encoding.
Benefits of trainable encodings for the expressivity of quantum models have been demonstrated numerically in [38, 39, 40] and theoretically in [16, 17]. In the following, we discuss the importance of trainable encodings for robustness and generalization. Recall that the Lipschitz bound (5), which we showed to be a crucial quantifier of robustness and generalization of quantum models, only depends on the observable $\mathcal{M}$ and on the data encoding $w_{j}$, $H_{j}$, but is independent of the parameters $\theta_{j}$. Hence, in the quantum model $f_{\varphi}^{\mathrm{f}}(x)$ with fixed encoding, the Lipschitz bound (5) cannot be influenced during training and, instead, is fixed a priori via the choice of the Hermitian generators $G_{j}$. As a result, training has a limited effect on robustness and generalization properties of fixed-encoding quantum models.

### v3

@@ Abstract
Adversarial robustness and generalization are both crucial properties of reliable machine learning models. In this paper, we study these properties in the context of quantum machine learning based on Lipschitz bounds. We derive parameter-dependent Lipschitz bounds for quantum models with trainable encoding, showing that the norm of the data encoding has a crucial impact on the robustness against data perturbations. Further, we derive a bound on the generalization error which explicitly involves the parameters of the data encoding. Our theoretical findings give rise to a practical strategy for training robust and generalizable quantum models by regularizing the Lipschitz bound in the cost. Further, we show that, for fixed and non-trainable encodings, as those frequently employed in quantum machine learning, the Lipschitz bound cannot be influenced by tuning the parameters. Thus, trainable encodings are crucial for systematically adapting robustness and generalization during training. The practical implications of our theoretical findings are illustrated with numerical results.
1 Introduction
Robustness of machine learning (ML) models is an increasingly important property, especially when operating on real-world data subject to perturbations. In practice, there are various possible sources of perturbations such as noisy data acquisition or adversarial attacks. The latter are tiny but carefully chosen manipulations of the data, and they can lead to dramatic misclassification in neural networks [1, 2]. As a result, much research has been devoted to better understanding and improving adversarial robustness [3, 4, 5]. It is well-known that robustness is closely connected to generalization [1, 2, 6, 7, 8], i.e., the ability of a model to extrapolate beyond the training data. Intuitively, if a model is robust then small input changes only cause small output changes, thus counteracting the risk of overfitting.
@@ A Lipschitz bound of a model $f$ is any $L>0$ satisfying
for all $x_{1},x_{2}\in\mathcal{D}\subseteq\mathbb{R}^{d}$. By definition, Lipschitz bounds quantify the worst-case output change that can be caused by data perturbations and, thus, they provide a useful measure of adversarial robustness. Therefore, they are a well-established tool for characterizing robustness and generalization properties of ML models [2, 6, 9, 10, 11, 12, 13, 14, 15]. Lipschitz bounds cannot only be used to better understand these two properties, but they also allow one to improve them by regularizing the Lipschitz bound during training [2, 6, 16, 17, 18].
Figure 1: Schematic illustration of the quantum model and training setup considered in this work for an exemplary Fashion MNIST data set [19]. The data, $x$, enter the quantum circuit via a trainable encoding, i.e., they are encoded into unitary operators $U_{j,\Theta_{j}}(x)$ via an affine function $w_{j}^{\top}x+\theta_{j}$ with trainable parameters $w_{j}$, $\theta_{j}$. During training, we minimize a cost function consisting of the empirical loss as well as an additional regularization term penalizing the norms of the parameters $w_{j}$. This regularization reduces the Lipschitz bound of the quantum model w.r.t. data perturbations. We show with both theoretical and numerical results that smaller values of $\lVert w_{j}\rVert$, as encouraged by the proposed regularization, lead to more robust and generalizable quantum models.
@@
In this paper, we study the interplay of robustness and generalization in quantum machine learning (QML). Variational quantum circuits are a well-studied class of quantum models [20, 21, 22, 23] and they promise benefits over classical ML in various aspects including trainability, expressivity, and generalization performance [24, 25]. Data re-uploading circuits generalize the classical variational circuits by concatenating a data encoding and a parametrized quantum circuit not only once but repeatedly, thus iterating between data- and parameter-dependent gates [26]. This alternation provides substantial improvements on expressivity, leading to a universal quantum classifier even in the single-qubit case [26, 27].
Just as in the classical case, robustness is crucial for quantum models. First, if QML is to provide benefits over classical ML, it is necessary to implement QML circuits which are robust w.r.t. quantum errors occurring due to imperfect hardware in the noisy intermediate-scale quantum (NISQ) era [28]. Questions of robustness of quantum models against such hardware errors have been studied, e.g., in [29, 30]. Lipschitz bounds can be used to study robustness of quantum algorithms against certain types of hardware errors, e.g., coherent control errors [31].
@@
However, robustness against hardware errors is entirely different from and independent of the robustness of a quantum model against data perturbations, which is the subject of this paper. The latter type of robustness has been studied in the context of quantum adversarial machine learning [32, 33]. Not surprisingly, just like their classical counterparts, quantum models are also vulnerable to adversarial attacks, both when operating based on classical data [34, 35] and quantum data [36, 34, 37, 38, 39, 40, 41]. To mitigate these attacks, it is desirable to design training schemes encouraging adversarial robustness of the resulting quantum model. Existing approaches in this direction include solving an (adversarial) min-max optimization problem during training [34] or adding adversarial examples to the training data set [42].
@@
Contribution
This paper presents a flexible and rigorous framework for robustness and generalization of quantum models, providing both a theoretical analysis as well as a simple regularization strategy which allows to systematically adapt robustness and generalization during training (see Figure 1 for an overview). More precisely, we first derive a Lipschitz bound of a given quantum model which explicitly involves the parameters of the data encoding. Based on this result, we propose a regularized training strategy penalizing the norm of the encoding parameters, which are considered trainable, in order to improve (adversarial) robustness of the model. Further, we derive a generalization bound which explicitly depends on the parameters of the quantum model and therefore does not share the limitations of existing uniform generalization bounds [49]. With numerical results, we demonstrate that the proposed Lipschitz bound regularization can indeed lead to substantial improvements in robustness and generalization of quantum models. Finally, given that the derived Lipschitz bound mainly depends on the norm of the data encoding, our results reveal the importance and benefits of trainable encodings over quantum circuits with a priori fixed encoding as frequently used in variational QML [20, 21, 22, 23, 24, 27].
Outline
The paper is structured as follows. In Section 2, we introduce the considered class of quantum models with trainable encodings and we state their Lipschitz bound. Next, in Section 3, we use the Lipschitz bound to study robustness of quantum models and to derive a regularization strategy for robust training whose benefits are demonstrated with numerical simulations. We then derive a generalization bound which depends explicitly on the data encoding parameters and we confirm this insight numerically by showing improved generalization under the proposed regularization strategy (Section 4). Further, in Section 5, we discuss an important implication of our results on the benefits of trainable encodings for robustness and generalization. Finally, Section 6 concludes the paper. In the appendix, we provide technical proofs, details on the numerical simulations, as well as additional theoretical and numerical results.
2 Quantum models and their Lipschitz bounds
We consider parametrized unitary operators of the form
@@ 2 Quantum models and their Lipschitz bounds
where $\Theta=\{\Theta_{j}\}_{j=1}^{N}$ comprises the set of trainable parameters. The quantum model considered in this paper consists of $U_{\Theta}(x)$ applied to |$\ket{0}$ and followed by a measurement w.r.t. the observable $\mathcal{M}$, i.e.,
@@
The formal derivation can be found in Appendix A. For a given set of parameters $w_{j}$, (5) allows to compute the Lipschitz bound of the corresponding quantum model. Note that $L_{\Theta}$ depends only on $w_{j}$ but it is independent of $\theta_{j}$. This fact plays an important role for potential benefits of trainable encodings since the parameters $w_{j}$ are not optimized during training for fixed-encoding circuits. We note that all results in this paper hold for arbitrary $p$-norms as long as the same $p$ is used for both vector and induced matrix norms.
3 Robustness of quantum models
Suppose we want to evaluate the quantum model $f_{\Theta}$ at $x$, i.e., we are interested in the value $f_{\Theta}(x)$, but we can only access $f_{\Theta}$ at some perturbed input $x^{\prime}=x+\varepsilon$ with an unknown $\varepsilon$. Such a setup can arise due to various reasons, e.g., $x$ may be the output of some physical process which can only be accessed via noisy sensors. The perturbation $\varepsilon$ may also be the result of an adversarial attack, i.e., a perturbation aiming to cause a misclassification by choosing $\varepsilon$ such that
$\lVert f_{\Theta}(x+\varepsilon)-f_{\Theta}(x)\rVert$ (6)
is maximized. In either case, to correctly classify $x$ despite the perturbation, we require that $f_{\Theta}(x+\varepsilon)$ is close to $f_{\Theta}(x)$, meaning that (6) is small. According to (1), a Lipschitz bound $L$ of $f_{\Theta}$ quantifies exactly this difference, implying that the maximum possible deviation of $f_{\Theta}(x+\varepsilon)$ from $f_{\Theta}(x)$ is bounded as
@@ 3 Robustness of quantum models
Regularizing the parameters $w_{j}$ encourages small norms of the data encoding and, thereby, small values of the Lipschitz bound $L_{\Theta}$. The hyperparameter $\lambda>0$ allows for a trade off between the two objectives of a small training loss and robustness/generalization in the cost function. Note that the regularization does not involve the $\theta_{j}$’s since they do not influence the Lipschitz bound (5), an issue we discuss in more detail in Section 5. Moreover, we do not introduce an explicit dependence of the regularization on $\mathcal{M}$ since we do not optimize over the observable in this paper. We note that penalty terms similar to the proposed regularization can be used for handling hard constraints in binary optimization via the quantum approximate optimization algorithm [50, 51]. Indeed, the above regularization can be interpreted as a penalty-based relaxation of the corresponding constrained training problem, i.e., of training a quantum model with Lipschitz bound below a specific value.
@@
We now evaluate our theoretical findings based on the circle classification problem from [52]: within a domain of $\mathcal{X}=[-1,+1]\times[-1,+1]$, a circle with radius $\sqrt{2/\pi}$ is drawn, and all data points inside the circle are labeled with $y=+1$, whereas points outside are labeled with $y=-1$, see Appendix C (Figure 5). For the quantum model with trainable encoding, we use general SU(2) operators and encode $w_{j}^{\top}x+\theta_{j}$ into the first two rotation angles. We repeat this encoding for each of the considered $3$ qubits, followed by nearest neighbour entangling gates based on CNOTs. Such a layer is then repeated $3$ times. As observable, we use $\mathcal{M}=Z\otimes Z\otimes Z$. The resulting circuit is illustrated in Appendix C (Figure 7). As norms in the regularized training problem (9), we employ $2$-norms, but we note that exploring different choices is an interesting future direction. For example, a $1$-norm regularization may enforce sparsity and, thereby, simplify implementations on NISQ hardware.
@@
The numerical results for the robustness simulations are shown in Figure 2, where we compare the worst-case test accuracy and Lipschitz bound of three trained models with different regularization parameter $\lambda\in\{0,0.2,0.5\}$. Additionally, the plot shows the accuracy of a quantum model with fixed encoding for the same numbers of qubits and layers (see (14) and Appendix C for details). The worst-case test accuracy of all models is obtained by sampling different noise samples $\varepsilon$ from $[-\bar{\varepsilon},+\bar{\varepsilon}]$ and, thereby, it approximates the accuracy resulting from an adversarial attack norm-bounded by $\bar{\varepsilon}$. As expected, all four models deteriorate with increasing noise level. For zero noise level $\bar{\varepsilon}=0$, the model with the largest regularization parameter $\lambda=0.5$ (and, hence, the smallest Lipschitz bound $L_{\Theta}=3.67$) has a smaller test accuracy than the non-regularized model with $\lambda=0$. This can be explained by a decrease in the training accuracy that is caused by the additional regularization in the cost. For increasing noise levels, however, the enhanced robustness outweighs the loss of training performance and, therefore, the model with $\lambda=0.5$ outperforms the model with $\lambda=0$. The fixed-encoding model achieves comparable performance to the trainable-encoding model with $\lambda=0.5$ for small noise and the worst performance among all models for high noise. These observations can be explained by the reduced expressivity but high Lipschitz bound of the fixed-encoding model. Finally, the model with $\lambda=0.2$ almost always outperforms the model with $\lambda=0$ and, in particular, it yields a higher test accuracy for small noise levels. This can be explained by the improved generalization performance caused by the regularization, an effect we discuss in more detail in the following.
4 Generalization of quantum models
@@
The Lipschitz bound (5) not only influences robustness but also has a crucial impact on generalization properties of the quantum model $f_{\Theta}$. Intuitively, a smaller Lipschitz bound implies a smaller variability of $f_{\Theta}$ and, therefore, reduces the risk of overfitting. This intuition is made formal via the following generalization bound.
@@
Figure 3: Results for the generalization simulations. The training setup is identical to the robustness simulations as described in Figure 2. As test set, we draw $10.000$ points uniformly at random and evaluate the trained models with different regularization parameter $\lambda$.
Theorem 4.1.
@@
(informal version) Consider a supervised learning setup with loss $\ell$ and data set $(x_{k},y_{k})\in\mathcal{X}\times\mathcal{Y}$ of size $n$ drawn according to the probability distribution $P$. For the quantum model $f_{\Theta}$ from (4), define the expected risk $\mathcal{R}(f_{\Theta})=\int_{\mathcal{X}\times\mathcal{Y}}\ell(y,f_{\Theta}(x))\mathrm{d}P(x,y)$ and the empirical risk $\mathcal{R}_{n}(f_{\Theta})=\frac{1}{n}\sum_{k=1}^{n}\ell(y_{k},f_{\Theta}(x_{k}))$. The generalization error of $f_{\Theta}$ is bounded as
$\Big{|}\mathcal{R}(f_{\Theta})-\mathcal{R}_{n}(f_{\Theta})\Big{|}$ (10)
$\leq$ $\,C_{1}\lVert\mathcal{M}\rVert\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert+\frac{C_{2}}{\sqrt{n}}$
@@
The detailed version and proof of Theorem 4.1 are provided in Appendix B. Generalization bounds as in (10) quantify the ability of $f_{\Theta}$ to generalize beyond the available data. The bound (10) depends on the data encoding via $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ and on the observable via $\lVert\mathcal{M}\rVert$. In particular, $f_{\Theta}$ achieves a small generalization error if its Lipschitz bound $L_{\Theta}$ is small and the size $n$ of the data set is large. Note, however, the following fundamental trade-off: A too small Lipschitz bound $L_{\Theta}$ may limit the expressivity of $f_{\Theta}$ and, therefore, lead to a high empirical risk $\mathcal{R}_{n}(f_{\Theta})$, in which case the generalization bound (10) is meaningless. In conclusion, Theorem 4.1 implies a small expected risk $\mathcal{R}(f_{\Theta})$ if $n$ is large, $L_{\Theta}$ is small, and $f_{\Theta}$ has a small empirical risk $\mathcal{R}_{n}(f_{\Theta})$. In contrast to existing generalization bounds [24, 45, 46, 47, 48], the bound (10) is not uniform and explicitly involves the Lipschitz bound (5), i.e., the parameters of the data encoding. Hence, Theorem 4.1 does not share the limitations of uniform QML generalization bounds [49].
@@
5 Benefits of trainable encodings
A popular class of quantum models is obtained by constructing circuits which alternate between data- and parameter-dependent gates, i.e., replacing $U_{\Theta}(x)$ in (3) by
$U^{\mathrm{f}}_{\phi}(x)=W(\phi_{L})V(x)\cdots W(\phi_{1})V(x),$ (11)
compare [20, 21, 22, 23, 24, 27]. The unitary operators $V,W$ are given by
$V(x)$ $=e^{-ix_{D}G_{D}}\cdots e^{-ix_{1}G_{1}},$ (12)
$W(\phi_{j})$ $=e^{-i\phi_{j,p}S_{p}}\cdots e^{-i\phi_{j,1}S_{1}}$ (13)
for trainable parameters $\phi_{j}$ and generators $G_{i}=G_{i}^{\dagger}$, $S_{i}=S_{i}^{\dagger}$. The corresponding quantum model is given by
$f^{\mathrm{f}}_{\phi}(x)=\braket{0}{U^{\mathrm{f}}_{\phi}(x)^{\dagger}\mathcal{M}U^{\mathrm{f}}_{\phi}(x)}{0},$ (14)
see Figure 4.
It is not hard to show that the parametrized quantum circuit $U_{\Theta}(x)$ in (3) generalizes the one in (11). Indeed, $U_{j,\Theta_{j}}(x)$ in (2) reduces to either
$e^{-ix_{j}G_{j}}\quad\text{or}\quad e^{-i\phi_{j}S_{j}}$ (15)
for suitable choices of $w_{j}$, $\theta_{j}$, and $H_{j}$. Note that the data encoding of the quantum model $f^{\mathrm{f}}_{\phi}(x)$ is fixed a priori via the choice of $w_{j}$ and, in particular, it cannot be influenced during training. Therefore, we refer to $f^{\mathrm{f}}_{\phi}(x)$ as a quantum model with fixed encoding, in contrast to $f_{\Theta}(x)$ in (4) which contains trainable parameters $w_{j}$ and, therefore, a trainable encoding.
Benefits of trainable encodings for the expressivity of quantum models have been demonstrated numerically in [53, 54, 55] and theoretically in [26, 27]. In the following, we discuss the importance of trainable encodings for robustness and generalization. Recall that the Lipschitz bound (5), which we showed to be a crucial quantifier of robustness and generalization of quantum models, only depends on the observable $\mathcal{M}$ and on the data encoding $w_{j}$, $H_{j}$, but is independent of the parameters $\theta_{j}$. Hence, in the quantum model $f_{\phi}^{\mathrm{f}}(x)$ with fixed encoding, the Lipschitz bound (5) cannot be influenced during training and, instead, is fixed a priori via the choice of the Hermitian generators $G_{j}$. As a result, training has a limited effect on robustness and generalization properties of fixed-encoding quantum models.